A charge of 28.0 nC is placed in a uniform electric field



that is directed vertically upward and has a magnitude of



4.00 104 V/m. What work is done by the electric force when



the charge moves (a) 0.450 m to the right; (b) 0.670 m upward;



(c) 2.60 m at an angle of 45.0° downward from the horizontal?

1
Expert's answer
2022-04-21T17:44:52-0400

"А=qEr\cos\alpha,"

1.

"\alpha=90°,\\ 
A=0~J,"

2.

"\alpha=0°,\\
A=750~\mu J,"

3.

"\alpha=-45°,\\
A=-2.06~mJ."